Chris has a change jar that contains $2.19 in quarters and pennies. He has 11 more pennies than quarters. How many of each type of coin does he have?
Tiana has a change jar that contains $3.50 in dimes and nickels. She has 12 times as many nickels as dimes. How many of each type of coin does she have?
Kelsey has a change jar that contains $2.45 in nickels and quarters. The number of nickels is 7 more than two times the number of quarters. How many of each type of coin does she have?

Found 4 solutions by Boreal, mccravyedwin, Alan3354, ikleyn:
Answer by Boreal(15235)

Here's the first one
quarters=x
pennies=x+11
.25x+.01(x+11)=2.19
.26x+.11=2.19
.26x=2.08
x=8 quarters
x+11=19 pennies

Answer by mccravyedwin(407)

Instead of doing your problems for you, I'll change only the numbers and work
problems exactly like yours step by step, that you can use as models to solve
your problems by.  Notice that the numbers are different but they are worked
exactly the same way:
Chris has a change jar that contains $2.51 in quarters and pennies. He
has 17 more pennies than quarters. How many of each type of coin does he have?
q = number of quarters 
$0.25q = value of all the quarters
q+17 = number of pennies   <-- the more...than... tells us to ADD q and 17.
$0.01(q+17) = value of all the pennies



$0.25q + $0.01(q+17) = $2.51
   25q +     1(q+17) =   251
       25q + 1q + 17 = 251
            26q + 17 = 251
Subtract 17 from both sides
                 26q = 234
Divide both sides by 26
                   q = 9     <-- So Chris has 9 quarters   

q+17 = number of pennies = 9+17 = 26 <-- So Chris has 26 pennies

--------------------------------------------------

Tiana has a change jar that contains $3.90 in dimes and nickels. She has
11 times as many nickels as dimes. How many of each type of coin does she have?

d = number of dimes 
$0.10d = value of all the dimes
11∙d = number of nickels    <--the word "times" tells us to multiply 11 by d.
$0.05(11∙d) = value of all the nickels



$0.10d + $0.05(11∙d) = $3.90
   10d +     5( 11d) =   390
           10d + 55d = 390
                 65d = 390
Divide both sides by 65
                   d = 6     <-- So Chris has 6 dimes   

11∙d = number of nickels = 11∙6 = 66 <-- So Tiana has 66 nickels

-------------------------------------------------------------------------

Kelsey has a change jar that contains $2.40 in nickels and quarters. The number
of nickels is 8 more than three times the number of quarters. How many of each
type of coin does she have?
q = number of quarters 
$0.25q = value of all the quarters
3q+8 = number of nickels  <--the "more than three times" tells us to ADD 3q and  7.
$0.05(3q+8) = value of all the nickels



$0.25q + $0.05(3q+8) = $2.40
   25q +     5(3q+8) =   240
      25q + 15q + 40 = 240
            40q + 40 = 240
Subtract 40 from both sides
                 40q = 240
Divide both sides by 35
                   q = 5     <-- So Kelsey has 5 quarters   

3q+8 = number of nickels = 3(5)+8 = 15+8 = 23 <-- So Kelsey has 23 nickels

Edwin


Answer by Alan3354(69443)

Chris has a change jar that contains $2.19 in quarters and pennies. He has 11 more pennies than quarters many of each type of coin does he have?
-------------
Set the "11 more pennies" aside.
Now it's sets of 26 cents each worth $2.08
208/26 = 8 sets
---> 8 Q's and 19 P's

Answer by ikleyn(52797)

.

To solve the first problem, you DO NOT NEED to write / (to solve) equations.

It is enough to look into this number $2.19 to see / (to guess) the answer MENTALLY :


                        19 pennies, and  19-11 = 8  quarters.


=============

Please NEVER pack more than one problem per post.